### v2

# Relation between generalized and ordinary cluster algebras

Abstract.
Recently, Ramos and Whiting showed that any generalized cluster algebra of geometric type is isomorphic to a quotient of a subalgebra of a certain cluster algebra. Based on their idea and method, we show that the same property holds for any generalized cluster algebra with $y$-variables in an arbitrary semifield. We also present the relations between the $C$-matrices, the $G$-matrices, and the $F$-polynomials of a generalized cluster pattern and those of the corresponding composite cluster pattern.
1. Introduction
Cluster algebras were introduced by Fomin and Zelevinsky [FZ02]. They naturally appear in several areas of mathematics and have been studied and developed by many authors. In [CS14], Chekhov and Shapiro introduced a natural generalization of cluster algebras called generalized cluster algebras (GCAs) by replacing the binomial exchange polynomials of (ordinary) cluster algebras (CAs) with the multinomial ones. It turned out that GCAs share several important properties with CAs, for example, the Laurent phenomenon [CS14], the separation formulas [NAK15b], quantizations [NAK15a, BCD+18, FPY24], scattering diagram formulation [CKM21, MOU24], the Laurent positivity [MOU24, BLM25b, BLM25a], etc.
Suggested by the question in [CS14, §5], Ramos and Whiting [RW25] recently showed that any GCA $\mathcal{A}^{\mathrm{g}}$ of geometric type (i.e., with coefficients in a tropical semifield) is isomorphic to a subquotient of a certain CA $\mathcal{A}$, that is, a quotient of a subalgebra of $\mathcal{A}$. This identification finally reveals the true nature of GCAs. Let us summarize the main idea in [RW25] without explaining details. Firstly, the exchange matrices of $\mathcal{A}$ are given by certain extensions (“enlargements”) of the exchange matrices of $\mathcal{A}^{\mathrm{g}}$ so that the “composite mutations” of the enlarged exchange matrices are compatible with the generalized mutations of the exchange matrices of $\mathcal{A}^{\mathrm{g}}$. Secondly, we consider the product of the cluster variables in $\mathcal{A}$ mutated by each composite mutation. We restrict our attention to the subalgebra $\mathcal{A}^{\mathrm{c}}$ of $\mathcal{A}$ generated by these cluster monomials. Lastly, we impose a certain relation on $\mathcal{A}^{\mathrm{c}}$ so that the exchange polynomials of $\mathcal{A}^{\mathrm{g}}$ coincide with the exchange polynomials of these cluster monomials in $\mathcal{A}^{\mathrm{c}}$. (This implies the factorization of the exchange polynomials of $\mathcal{A}^{\mathrm{g}}$ into binomials under the relation.) Then, $\mathcal{A}^{\mathrm{g}}$ is isomorphic to the quotient of $\mathcal{A}^{\mathrm{c}}$ by the relation ideal.
Due to the importance of the above result, in this paper we apply the same idea and method to GCAs with coefficients in an arbitrary semifield and obtain a parallel result for them. To do it, in contrast to [RW25], we concentrate on the class of GCAs with $y$-variables in [NAK15b, NR16, NAK24], which is parallel to the one of CAs in [FZ07]. Compared with the formulation of GCAs in [RW25], the advantage of ours is two-folded. Firstly, we are away from the specific problem arising from the exchange matrix presentation of tropical coefficients in [RW25]. (A considerable part of the paper [RW25] is devoted to resolve the issues related to it.) Secondly, the parallelism of the factorization mechanism between cluster variables ($x$-variables) and coefficients ($y$-variables) is transparent. We also present the relations between the $C$-matrices, the $G$-matrices, and the $F$-polynomials of a generalized cluster pattern and those of the corresponding composite cluster pattern.
Though mostly our method is a straightforward extension of the one in [RW25], there are some subtleties which we need to take care of. As explained above, to obtain the desired factorization of the exchange polynomials for the $x$-variables of a GCA $\mathcal{A}^{\mathrm{g}}$, we need to impose a certain relation on the subalgebra $\mathcal{A}^{\mathrm{c}}$ of the corresponding CA $\mathcal{A}$. The relation ideal of $\mathcal{A}^{\mathrm{c}}$ is prime; thus, its quotient is an integral domain. (This is consistent with the fact that the quotient is isomorphic to $\mathcal{A}^{\mathrm{g}}$, which is an integral domain.) However, if we extend the relation to $\mathcal{A}$, the relation ideal of $\mathcal{A}$ might be non-prime so that its quotient is not an integral domain, in general. A similar situation occurs also for $y$-variables, and we need to handle it carefully. In fact, clarifying these details is one of the main achievement of the paper. The summary of this point is given in the diagram (5.10).
The content of the paper is as follows. In Section 2, we recall the definition of generalized cluster algebras with $y$-variables. In Section 3, we introduce composite mutations and composite cluster patterns in some ordinary cluster algebras. In Section 4, we realize the generalized $Y$-patterns by the composite ones. In Section 5, we realize the generalized cluster patterns and the generalized cluster algebras by the composite ones. In Section 6, we present the relations between the $C$-matrices, the $G$-matrices and the $F$-polynomials of generalized cluster algebras and those of composite ones.
We thank Zhichao Chen and Yasuaki Gyoda for discussions. We thank Woojin Choi for pointing out the error in the early version of the manuscript. This work is supported in part by JSPS grant No. JP22H01114 and No. JP25KJ1438. Ryota Akagi is also supported by Chubei Itoh Foundation.
2. Generalized cluster algebras with $y$-variables
In this section, we recall the definition of generalized cluster algebras (GCAs) with $y$-variables following [NAK15b, NR16, NAK24].
2.1. Semifields
Following [FZ02], we recall the definition of a semifield.
Definition 2.1 (Semifield).
A triplet $(\mathbb{P},\oplus,\cdot)$ is called a semifield if it satisfies the following conditions.
  • $\mathbb{P}$ is an abelian group with respect to the multiplication $\cdot$.
  • $\mathbb{P}$ is a commutative semigroup with respect to the addition $\oplus$.
  • The distributive property $p(q_{1}\oplus q_{2})=pq_{1}\oplus pq_{2}$ holds for any $p,q_{1},q_{2}\in\mathbb{P}$.
For any semifield $\mathbb{P}$, we set $\mathbb{P}_{0}=\mathbb{P}\sqcup\{0\}$, and we extend the addition and the multiplication to $\mathbb{P}_{0}$ by $0\oplus p=p\oplus 0=p$ and $0\cdot p=p\cdot 0=0$ for any $p\in\mathbb{P}_{0}$. Since $\mathbb{P}_{0}$ is not an abelian group with respect to $\cdot$, it is not a semifield in the above sense.
Let $\mathbb{ZP}$ be the group ring of $(\mathbb{P},\cdot)$ over $\mathbb{Z}$. We set
(2.1) $\mathbb{NP}=\left\{\left.\ \sum_{j=1}^{r}a_{j}p_{j}\in\mathbb{ZP}\ \right|\ r\in\mathbb{Z}_{\geq 0},\ a_{j}\in\mathbb{Z}_{\geq 0},\ p_{j}\in\mathbb{P}\right\},$
and we define an additive and multiplicative homomorphism $\pi_{\mathbb{P}_{0}}:\mathbb{NP}\rightarrow\mathbb{P}_{0}$ by
(2.2) $\pi_{\mathbb{P}_{0}}\left(\sum_{j=1}^{r}a_{j}p_{j}\right)=\bigoplus_{j=1}^{r}a_{j\oplus}p_{j}\quad(a_{j}\in\mathbb{Z}_{\geq 0},p_{j}\in\mathbb{P}),$
where $a_{\oplus}p=p\oplus p\oplus\cdots\oplus p$ ($a\in\mathbb{Z}_{\geq 0}$, $p\in\mathbb{P}$) is the sum of $a$ terms of $p$ with respect to the addition $\oplus$ in $\mathbb{P}$. This should be distinguished from $ap=p+p+\cdots+p$, which is the sum in $\mathbb{ZP}$.
The ring $\mathbb{ZP}$ is an integral domain [FZ02]. The fraction field of $\mathbb{ZP}$ is denoted by $\mathbb{QP}$.
2.2. Generalized seeds and GCAs
When we consider a GCA with $y$-variables, we always fix the following data:
  • a positive integer $n\in\mathbb{Z}_{\geq 1}$ called the rank.
  • an $n$-tuple of positive integers ${\bf r}=(r_{1},\dots,r_{n})\in\mathbb{Z}_{\geq 1}^{n}$ called the mutation degree.
  • a semifield $\mathbb{P}$ called the coefficient semifield.
  • a field of rational functions $\mathcal{F}$ in $n$-variables with coefficients in $\mathbb{QP}$ called the ambient field.
Definition 2.2.
A quadruplet $\Sigma=({\bf x},{\bf y},{\bf Z},B)$ is called a generalized seed if it satisfies the following conditions.
  • ${\bf x}=(x_{1},\dots,x_{n})$ is an $n$-tuple of elements in $\mathcal{F}$ forming a free generating set. Each $x_{i}$ is called a cluster variable ($x$-variable).
  • ${\bf y}=(y_{1},\dots,y_{n})$ is an $n$-tuple of elements in $\mathbb{P}$. Each $y_{i}$ is called a coefficient ($y$-variable).
  • ${\bf Z}=(Z_{1}(u),\dots,Z_{n}(u))$ is an $n$-tuple of polynomials $Z_{i}(u)\in\mathbb{NP}[u]$ of the form
    (2.3) $Z_{i}(u)=1+\hat{z}_{i1}u+\hat{z}_{i2}u^{2}+\cdots+\hat{z}_{i,r_{i}-1}u^{r_{i}-1}+u^{r_{i}}\quad(\hat{z}_{il}\in\mathbb{NP}).$
    Each $Z_{i}(u)$ is called an exchange polynomial. We allow some of $\hat{z}_{il}$ to be zero. We set $\hat{z}_{i0}=\hat{z}_{ir_{i}}=1$.
  • $B=(b_{ij})\in\mathrm{Mat}_{n}(\mathbb{Z})$ is a skew-symmetrizable matrix, that is, there is a rational diagonal matrix $D=\mathrm{diag}(d_{1},\dots,d_{n})$ with $d_{i}>0$ such that $DB$ is skew-symmetric. It is called an exchange matrix.
For a generalized seed $\Sigma$, we define $\hat{y}-$variables as
(2.4) $\hat{y}_{k}=y_{k}\prod_{j=1}^{n}x_{j}^{{b}_{jk}}\quad(k=1,\dots,n).$
For a generalized seed $({\bf x},{\bf y},{\bf Z},B)$, we sometimes focus on the $y$-variables $\bf y$} only. In this case, the triplet $({\bf y},{\bf Z},B)$ is called a generalized $Y$-seed.
For an exchange polynomial $Z_{k}(u)=\sum_{l=0}^{r_{k}}\hat{z}_{kl}u^{l}$ and $y\in\mathbb{P}$, we define the specialization $Z_{k}|_{\mathbb{P}_{0}}(y)$ of $Z_{k}(u)$ at $y$ by
(2.5) $Z_{k}|_{\mathbb{P}_{0}}(y)=\bigoplus_{l=0}^{r_{k}}\pi_{\mathbb{P}_{0}}(\hat{z}_{kl})y^{l}.$
Since $\hat{z}_{k0}=\hat{z}_{kr_{k}}=1$, we have $Z_{k}|_{\mathbb{P}_{0}}(y)\in\mathbb{P}$. We write $\pi_{\mathbb{P}_{0}}(\hat{z}_{kl})\in\mathbb{P}_{0}$ as $z_{kl}$, for simplicity.
Definition 2.3.
Let $\Sigma=({\bf x},{\bf y},{\bf Z},B)$ be a generalized seed. Then, the mutation $\mu_{k}^{\bf r}(\Sigma)=({\bf x}^{\prime},{\bf y}^{\prime},{\bf Z}^{\prime},B^{\prime})$ in direction $k=1,2,\dots,n$ with the mutation degree $\bf r$} is defined as follows:
  • ${\bf x}^{\prime}=(x_{1},\dots,x_{k}^{\prime},\dots,x_{n})$, where
    (2.6) $x_{k}^{\prime}=x_{k}^{-1}\left(\prod_{j=1}^{n}x_{j}^{[-b_{jk}]_{+}}\right)^{r_{k}}\frac{Z_{k}(\hat{y}_{k})}{Z_{k}|_{\mathbb{P}_{0}}(y_{k})}.$
  • ${\bf y}^{\prime}=(y_{1}^{\prime},\dots,y_{n}^{\prime})$, where
    (2.7) $y_{i}^{\prime}=\begin{cases}y_{k}^{-1}&i=k,\\ y_{i}\left(y_{k}^{[b_{ki}]_{+}}\right)^{r_{k}}Z_{k}|_{\mathbb{P}_{0}}(y_{k})^{-b_{ki}}&i\neq k.\end{cases}$
  • ${\bf Z}^{\prime}=(Z_{1}(u),\dots,Z_{k}^{\mathrm{rec}}(u),\cdots,Z_{n}(u))$, where $Z_{k}^{\mathrm{rec}}(u)=u^{r_{k}}Z_{k}(u^{-1})$ is the reciprocal polynomial of $Z_{k}(u)$.
  • $B^{\prime}=(b^{\prime}_{ij})=\mu^{\bf r}_{k}(B)$, where
    (2.8) $b^{\prime}_{ij}=\begin{cases}-b_{ij}&i=k\ \textup{or}\ j=k,\\ b_{ij}+r_{k}(b_{ik}[b_{kj}]_{+}+[-b_{ik}]_{+}b_{kj})&i,j\neq k.\end{cases}$
Let $\mathbb{T}_{n}$ be the $n$-regular tree with edges labeled by $1,2,\dots,n$. We say that a pair of vertices $t,t^{\prime}\in\mathbb{T}_{n}$ are $k$-adjacent ($k=1,\dots,n$) if $t$ and $t^{\prime}$ are connected with an edge labeled by $k$. Let us fix a given vertex (the initial vertex) $t_{0}\in\mathbb{T}_{n}$. Then, the collection of generalized seeds ${\bf\Sigma}=\{\Sigma_{t}\}_{t\in\mathbb{T}_{n}}$ is uniquely determined by the initial seed $\Sigma_{t_{0}}$ and the condition $\Sigma_{t^{\prime}}=\mu_{k}^{\bf r}(\Sigma_{t})$ for any pair of $k$-adjacent vertices $t,t^{\prime}\in\mathbb{T}_{n}$. The collection $\bf\Sigma$} is called the generalized cluster pattern with the initial seed $\Sigma_{t_{0}}$. The collection of the generalized $Y$-seeds ${\bf\Upsilon}=\{({\bf y}_{t},{\bf Z}_{t},B_{t})\}_{t\in\mathbb{T}_{n}}$ of $\bf\Sigma$} is called the generalized $Y$-pattern of $\bf\Sigma$}.
Definition 2.4 (Generalized cluster algebra).
Let $\bf\Sigma$} be a generalized cluster pattern. Let $\mathcal{X}$ be the set of all cluster variables appearing in $\bf\Sigma$}. Then, the generalized cluster algebra $\mathcal{A}({\bf\Sigma})$ of $\bf\Sigma$} is defined by the $\mathbb{ZP}$-subalgebra $\mathbb{ZP}[\mathcal{X}]$ of $\mathcal{F}$ generated by $\mathcal{X}$.
When ${\bf r}=(1,1,\dots,1)$, a generalized cluster pattern and a GCA reduce to an ordinary cluster pattern and an ordinary cluster algebra (CA) in [FZ07]. In this case, its exchange polynomials are given by $Z_{i}(u)=1+u$. So, we omit $\bf Z$} and write an ordinary seed as $\Sigma=({\bf x},{\bf y},B)$. The mutation $\mu_{k}^{\bf r}$ also reduces to the ordinary mutation $\mu_{k}$ in [FZ07].
3. Composite cluster patterns
In this section, following [RW25], for a generalized cluster pattern, we introduce a corresponding composite cluster pattern whose mutations are given by the composite mutations.
3.1. Enlargements of exchange matrices
For the mutation degree ${\bf r}=(r_{1},\dots,r_{n})$ of a generalized seed $\Sigma$, the sum $\mathcal{N}=r_{1}+\cdots+r_{n}$ is called the pseudo-rank of $\Sigma$, which will be the rank of a corresponding ordinary seed. Let $\mathbbm{1}^{l\times m}$ be the $l\times m$ matrix whose all components are $1$.
Definition 3.1 (Enlargement).
Let $B$ be the exchange matrix of a given generalized seed with mutation degree $\bf r$}. We define an $\mathcal{N}\times\mathcal{N}$ matrix $\mathcal{B}\in\mathrm{Mat}_{\mathcal{N}}(\mathbb{Z})$ with the $n\times n$ block decomposition whose $(i,j)$th block $\mathcal{B}^{ij}$ is given by
(3.1) $\mathcal{B}^{ij}=b_{ij}\mathbbm{1}^{r_{i}\times r_{j}}.$
We call $\mathcal{B}$ the enlargement of $B$ with respect to $\bf r$}. Let $\mathcal{B}^{ij}_{lm}$ denote the $(l,m)$th entry of the $(i,j)$th block $\mathcal{B}^{ij}$.
Example 3.2.
When ${\bf r}=(2,3)$ and $B=\left(\begin{smallmatrix}0&1\\ -2&0\end{smallmatrix}\right)$, the enlargement of $B$ is given by
(3.2) $\mathcal{B}=\left(\begin{array}[]{cc|ccc}0&0&1&1&1\\ 0&0&1&1&1\\ \hline\cr-2&-2&0&0&0\\ -2&-2&0&0&0\\ -2&-2&0&0&0\end{array}\right).$
Let us define the index set
(3.3) $\mathcal{D}=\{(i,l)\mid i=1,\dots,n;\ l=1,\dots,r_{i}\}.$
We write $\mathcal{B}^{ij}_{lm}$ also as $\mathcal{B}_{il,jm}$ ($(i,l),(j,m)\in\mathcal{D}$). We consider an ordinary seed whose exchange matrix is $\mathcal{B}$. Then, for each $(i,l)\in\mathcal{D}$, the mutation in direction $r_{1}+r_{2}+\cdots+r_{i-1}+l$ is denoted by $\mu_{il}$ in accordance with the above parametrization of $\mathcal{B}$.
Definition 3.3 (Composite mutation).
For each $k=1,2,\dots,n$, we define the composite mutation $\mu^{\mathrm{c}}_{k}$ for an ordinary seed $({\bf x},{\bf y},\mathcal{B})$ by
(3.4) $\mu^{\mathrm{c}}_{k}=\mu_{kr_{k}}\circ\cdots\circ\mu_{k2}\circ\mu_{k1}.$
The composite mutation $\mu^{\mathrm{c}}_{k}$ is independent of the order of the composition thanks to $\mathcal{B}^{kk}=O$. Moreover, the composite mutations are compatible with the enlargement as follows.
Lemma 3.4 ([RW25, Lem. 22]).
For a generalized cluster pattern with mutation degree $\bf r$}, let $\{B_{t}\}_{t\in\mathbb{T}_{n}}$ be the collection of its exchange matrices. Let $\mathcal{B}_{t}$ be the enlargement of $B_{t}$ with $\bf r$}. Then, for any pair of $k$-adjacent vertices $t,t^{\prime}\in\mathbb{T}_{n}$, we have
(3.5) $\mathcal{B}_{t^{\prime}}=\mu^{\mathrm{c}}_{k}(\mathcal{B}_{t}).$
Proof.
Let $\mathcal{B}$ be the enlargement of $B=(b_{ij})$. Let $B^{\prime}=(b_{ij}^{\prime})=\mu_{k}^{\bf r}(B)$ and $\mathcal{B}^{\prime}=\mu_{k}^{\mathrm{c}}(\mathcal{B})$. Then, for $(i,l),(j,m)\in\mathcal{D}$ with $i,j\neq k$, we have
(3.6) $(\mathcal{B}^{\prime})^{ij}_{lm}$ $=\mathcal{B}^{ij}_{lm}+\sum_{p=1}^{r_{k}}(\mathcal{B}^{ik}_{lp}[\mathcal{B}^{kj}_{pm}]_{+}+[-\mathcal{B}^{ik}_{lp}]_{+}\mathcal{B}^{kj}_{pm})$
$=b_{ij}+r_{k}(b_{ik}[b_{kj}]_{+}+[-b_{ik}]_{+}b_{kj})=b^{\prime}_{ij}.$
For $i=k$ or $j=k$, we have $(\mathcal{B}^{\prime})^{ij}_{lm}=-\mathcal{B}^{ij}_{lm}=-b_{ij}=b_{ij}^{\prime}$ because $\mathcal{B}^{kk}=O$. ∎
3.2. Composite cluster patterns
Let $\mathcal{B}$ be the matrix in (3.1). In accordance with the parametrization of $\mathcal{B}$, when we consider an ordinary seed $\Sigma=({\bf x},{\bf y},\mathcal{B})$, we parametrize $\bf x$} and $\bf y$} by $\mathcal{D}$ as ${\bf x}=(x_{il})_{(i,l)\in\mathcal{D}}$ and ${\bf y}=(y_{il})_{(i,l)\in\mathcal{D}}$. Then, we apply the composite mutations $\mu^{\mathrm{c}}_{k}$ in (3.4) to $\Sigma$. Explicitly, we have the following formula.
Lemma 3.5.
For an ordinary seed $\Sigma=({\bf x},{\bf y},\mathcal{B})$, the composite mutation $\mu^{\mathrm{c}}_{k}(\Sigma)=({\bf x}^{\prime},{\bf y}^{\prime},\mu^{\mathrm{c}}_{k}(\mathcal{B}))$ in direction $k=1,\dots,n$ is given as follows:
(3.7) $x^{\prime}_{il}$ $=\begin{cases}\displaystyle{x_{kl}^{-1}\biggl(\prod_{(j,m)\in\mathcal{D}}x_{jm}^{[-{b}_{jk}]_{+}}\biggr)\frac{1+\hat{y}_{kl}}{1\oplus y_{kl}}}&i=k,\\ x_{il}&i\neq k.\end{cases}$
(3.8) $y^{\prime}_{il}$ $=\begin{cases}y_{kl}^{-1}&i=k,\\ \displaystyle{y_{il}\left(\prod_{m=1}^{r_{k}}y_{km}\right)^{[{b}_{ki}]_{+}}\left(\prod_{m=1}^{r_{k}}(1\oplus y_{km})\right)^{-{b}_{ki}}}&i\neq k.\end{cases}$
Proof.
They follow from $\mathcal{B}^{ij}_{lm}=b_{ij}$, (2.6), and (2.7) with ${\bf r}=(1,\dots,1)$. ∎
Definition 3.6 (Composite cluster pattern).
In the definition of an ordinary cluster pattern, we replace the mutations $\mu_{k}$ ($k=1,\dots,\mathcal{N}$) with the composite ones $\mu^{\mathrm{c}}_{k}$ ($k=1,\dots,n$). We call the resulting collection ${\bf\Sigma}^{\mathrm{c}}=\{\Sigma^{\mathrm{c}}_{t}\}_{t\in\mathbb{T}_{n}}$ of ordinary seeds the composite cluster pattern with the initial seed $\Sigma_{t_{0}}^{\mathrm{c}}$. This is a subcollection of the ordinary cluster pattern $\bf\Sigma$} of rank $\mathcal{N}$ with the same initial seed $\Sigma_{t_{0}}=\Sigma_{t_{0}}^{\mathrm{c}}$. We also define the composite $Y$-pattern ${\bf\Upsilon}^{\mathrm{c}}$ of ${\bf\Sigma}^{\mathrm{c}}$ as the collection of the $Y$-seeds in ${\bf\Sigma}^{\mathrm{c}}$.
4. Realization of generalized $Y$-pattern
In this section, for a given generalized $Y$-pattern ${\bf\Upsilon}^{\mathrm{g}}$, we construct a composite $Y$-pattern ${\bf\Upsilon}^{\mathrm{c}}$ such that ${\bf\Upsilon}^{\mathrm{g}}$ is recovered from ${\bf\Upsilon}^{\mathrm{c}}$.
4.1. Extensions of semifields
We fix a generalized seed $\Sigma=({\bf x},{\bf y},{\bf Z},B)$ with coefficients in $\mathbb{P}$ and mutation degree $\bf r$}. Let us introduce formal variables $s_{il}$ ($(i,l)\in\mathcal{D}$) and set ${\bf s}=\{s_{il}\mid(i,l)\in\mathcal{D}\}$. Since $\mathbb{ZP}$ is an integral domain, the polynomial ring $\mathbb{ZP}[{\bf s}]$ in $\bf s$} is also so. Thus, we have the fraction field $\mathbb{QP}({\bf s})$ of $\mathbb{ZP}[{\bf s}]$. Let $\mathbb{NP}[{\bf s}]\subset\mathbb{ZP}[{\bf s}]$ be the set of all polynomials in $\bf s$} such that all coefficients belong to $\mathbb{NP}$ in (2.1). We introduce the universal semifield $\mathbb{QP}_{\mathrm{sf}}({\bf s})$ with coefficients in $\mathbb{P}$ by
(4.1) $\mathbb{QP}_{\mathrm{sf}}({\bf s})=\left\{\left.\frac{f}{g}\in\mathbb{QP}({\bf s})\ \right|f,g\in\mathbb{NP}[{\bf s}]\setminus\{0\}\right\}.$
This is a semifield with respect to the addition and multiplication of fractions in $\mathbb{QP}({\bf s})$. Below, we write $\mathbb{QP}_{\mathrm{sf}}({\bf s})$ as $\overline{\mathbb{P}}$, for simplicity.
Note that the addition in $\overline{\mathbb{P}}$ is different from the one in $\mathbb{P}$. When we consider cluster algebras in Section 5, we need to further consider the group rings $\mathbb{ZP}$ and $\mathbb{Z}\overline{\mathbb{P}}$, whose additions are also different. To distinguish them, we write these additions as follows:
  • $\oplus$ is the addition in $\mathbb{P}$.
  • $\bar{\oplus}$ is the addition in $\overline{\mathbb{P}}$.
  • ++ is the addition in $\mathbb{ZP}$ and $\mathbb{Z}\overline{\mathbb{P}}$.
Note that $\mathbb{NP}$ is a subset of both $\overline{\mathbb{P}}$ and $\mathbb{ZP}$. Since the additions $\bar{\oplus}$ and ++ coincide in $\mathbb{NP}$, we use both symbols depending on the context.
We would like to impose an equivalence relation $\sim$ on $\overline{\mathbb{P}}_{0}$, which is compatible with the addition and the multiplication, so that the following relation holds:
(4.2) $Z_{i}|_{\overline{\mathbb{P}}_{0}}(y)\sim(1\bar{\oplus}s_{i1}y)(1\bar{\oplus}s_{i2}y)\cdots(1\bar{\oplus}s_{ir_{i}}y).$
However, imposing such a relation directly on $\overline{\mathbb{P}}_{0}$ causes possible problems as follows.
(1) There is no guarantee that the resulting relation satisfies the primeness condition, i.e., $r_{1}r_{2}\sim 0$ implies $r_{1}\sim 0$ or $r_{2}\sim 0$. And, if the primeness condition does not hold, we have $1\sim 0$, so that the quotient $\overline{\mathbb{P}}_{0}/{\sim}$ collapses to $\{0\}$.
(2) Even if the primeness condition holds, there is no guarantee that the map $\mathbb{P}\rightarrow\overline{\mathbb{P}}_{0}/{\sim}$, $p\mapsto[p]$ is injective, in general.
To avoid these problems, for each $i=1,\dots,n$, we introduce the elementary symmetric polynomial in $s_{i1}$$s_{ir_{i}}$ of degree $l=1,\dots,r_{i}$ as
(4.3) $e_{il}=\mathop{\overline{\bigoplus}}_{1\leq m_{1}<m_{2}<\cdots<m_{l}\leq r_{i}}s_{im_{1}}s_{im_{2}}\cdots s_{im_{l}}\in\overline{\mathbb{P}}.$
Let ${\bf e}=\{{e}_{il}\mid(i,l)\in\mathcal{D}\}$. Let $\bf Z$} be the exchange polynomials of $\Sigma$. By extending the homomorphism $\pi_{\mathbb{P}_{0}}$ in (2.2), we define an additive and multiplicative homomorphism $\psi=\psi_{\bf Z}:\mathbb{NP}[{\bf e}]\to\mathbb{P}_{0}$ by
(4.4) $\psi(e_{il})=z_{il}.$
In particular, since $z_{ir_{i}}=\pi_{\mathbb{P}_{0}}(\hat{z}_{ir_{i}})=1$, we have
(4.5) $\psi(e_{ir_{i}})=\psi(s_{i1}s_{i2}\cdots s_{ir_{i}})=1\quad(i=1,\dots,n).$
We introduce the following subsemifield of $\overline{\mathbb{P}}$:
(4.6) $\mathbb{S}=\left\{\left.\frac{f}{g}\in\overline{\mathbb{P}}\ \right|\ f,g\in\mathbb{NP}[{\bf e}],\ \psi(f),\psi(g)\neq 0\right\}.$
We view $\mathbb{P}\subset\mathbb{S}$ by the natural identification $p=p/1$ ($p\in\mathbb{P}$). We obtain the following semifield homomorphism (denoted by the same symbol $\psi$)
(4.7) $\begin{array}[]{rccccc}\psi&:&\mathbb{S}&\to&\mathbb{P}&\\ &&\displaystyle{\frac{f}{g}}&\mapsto&\displaystyle{\frac{\psi(f)}{\psi(g)}}&(f,g\in\mathbb{NP}[{\bf e}]).\end{array}$
The map $\psi$ induces the equivalence relation in $\mathbb{S}$ such that
(4.8) $f\sim g\Longleftrightarrow\psi(f)=\psi(g)\quad(f,g\in\mathbb{S}).$
Let $[f]$ be the equivalence class of $f\in\mathbb{S}$, and let $\mathbb{S}/{\sim}$ be the quotient of $\mathbb{S}$ by $\sim$.
Lemma 4.1.
The quotient $\mathbb{S}/{\sim}$ is a semifield. Moreover, the map $\mathbb{P}\to\mathbb{S}/{\sim}$ defined by $p\mapsto[p]$ is a semifield isomorphism.
Proof.
The composition $\mathbb{P}\hookrightarrow\mathbb{S}\overset{\psi}{\to}\mathbb{P}$ is the identity map, where $\mathbb{P}\hookrightarrow\mathbb{S}$ is the canonical inclusion. Thus, $\psi$ is surjective. Then, by the fundamental theorem on homomorphisms, the claim holds. ∎
Now, we have the desired relation (4.2) in $\mathbb{S}$.
Lemma 4.2.
For any $y\in\mathbb{S}$ and $k=1,2,\dots,n$, the products $\prod_{l=1}^{r_{k}}(1\bar{\oplus}s_{kl}y)$ and $\prod_{l=1}^{r_{k}}(1\bar{\oplus}s_{kl}^{-1}y)$ belong to $\mathbb{S}$. Moreover, we have
(4.9) $\prod_{l=1}^{r_{k}}(1\bar{\oplus}s_{kl}y)\sim Z_{k}|_{\mathbb{S}_{0}}(y),\quad\prod_{l=1}^{r_{k}}(1\bar{\oplus}s_{kl}^{-1}y)\sim Z_{k}^{\mathrm{rec}}|_{\mathbb{S}_{0}}(y).$
Proof.
Let us prove the first relation. Since $\prod_{l=1}^{r_{k}}(1\bar{\oplus}s_{kl}y)=1\bar{\oplus}\mathop{\overline{\bigoplus}}_{l=1}^{r_{k}}e_{kl}y^{l}$, we have
(4.10) $\psi\left(\prod_{l=1}^{r_{k}}(1\bar{\oplus}s_{kl}y)\right)=1\oplus\bigoplus_{l=1}^{r_{k}}z_{kl}\psi(y)^{l}.$
The right hand side has a term $1\in\mathbb{P}$. So, it is nonzero. Thus, the product $\prod_{l=1}^{r_{k}}(1\bar{\oplus}s_{kl}y)$ belongs to $\mathbb{S}$. The first relation also follows from (4.10). The second relation is obtained from the first one by replacing $y$ with $y^{-1}$ and multiplying $y^{r_{k}}e_{kr_{k}}^{-1}\sim y^{r_{k}}$. ∎
4.2. Realization of generalized $Y$-patterns
From now on, we consider two kinds of patterns. The first one is the generalized cluster pattern ${\bf\Sigma}^{\mathrm{g}}=\{({\bf x}_{t}^{\mathrm{g}},{\bf y}_{t}^{\mathrm{g}},{\bf Z}_{t},B_{t})\}_{t\in\mathbb{T}_{n}}$, where the initial seed $\Sigma_{t_{0}}^{\mathrm{g}}$ is given by $\Sigma$ in the previous section. The second one is a composite cluster pattern ${\bf\Sigma}^{\mathrm{c}}=\{({\bf x}_{t}^{\mathrm{c}},{\bf y}_{t}^{\mathrm{c}},\mathcal{B}_{t})\}_{t\in\mathbb{T}_{n}}$ whose detail will be specified below.
Here, we focus on the generalized $Y$-pattern ${\bf\Upsilon}^{\mathrm{g}}$ of ${\bf\Sigma}^{\mathrm{g}}$ with the initial generalized $Y$-seed $({\bf y}_{t_{0}}^{\mathrm{g}},{\bf Z}_{t_{0}},B_{t_{0}})$. We introduce the corresponding initial ordinary $Y$-seed $({\bf y}_{t_{0}}^{\mathrm{c}},\mathcal{B}_{t_{0}})$ of rank $\mathcal{N}$ with coefficients in $\overline{\mathbb{P}}$, where $\mathcal{B}_{t_{0}}$ is the enlargement of $B_{t_{0}}$ in (3.1), and the $y$-variables are given by
(4.11) $y_{il;t_{0}}^{\mathrm{c}}=s_{il}y_{i;t_{0}}^{\mathrm{g}}\quad((i,l)\in\mathcal{D}).$
Then, we have the composite $Y$-pattern ${\bf\Upsilon}^{\mathrm{c}}$ with the initial $Y$-seed $({\bf y}_{t_{0}}^{\mathrm{c}},\mathcal{B}_{t_{0}})$.
For each $i=1,\dots,n$ and $t\in\mathbb{T}_{n}$, we recursively define the sign $\sigma_{i;t}\in\{\pm 1\}$ as follows: Set $\sigma_{i;t_{0}}=1$ for any $i=1,\dots,n$. For any pair of $k$-adjacent vertices $t,t^{\prime}\in\mathbb{T}_{n}$, let
(4.12) $\sigma_{k;t^{\prime}}=-\sigma_{k;t},\quad\sigma_{i;t^{\prime}}=\sigma_{i;t}\quad(i\neq k).$
Note that we have $Z_{i;t}(u)=Z_{i;t_{0}}(u)$ if $\sigma_{i;t}=1$, and $Z_{i;t}(u)=Z_{i;t_{0}}^{\mathrm{rec}}(u)$ if $\sigma_{i;t}=-1$. Then, by Lemma 4.2, we have
(4.13) $\prod_{l=1}^{r_{k}}(1\bar{\oplus}s_{kl}^{\sigma_{k;t}}y)\sim Z_{k;t}|_{\mathbb{S}_{0}}(y)\quad(y\in\mathbb{S}).$
Let $\mathbb{S}_{1}=\psi^{-1}(1)$, where $\psi$ is the one in (4.7). Then, the following fact holds.
Lemma 4.3 (cf. [RW25, Lem. 24]).
For any $(i,l)\in\mathcal{D}$ and $t\in\mathbb{T}_{n}$, we have
(4.14) $y_{il;t}^{\mathrm{c}}=s_{il}^{\sigma_{i;t}}y_{i;t}^{\mathrm{g}}\mathfrak{u}_{i;t},$
where $\mathfrak{u}_{i;t}\in\mathbb{S}_{1}$ is independent of $l=1,\dots,r_{i}$.
Proof.
When $t=t_{0}$, the claim follows from (4.11). Suppose that the claim holds for some $t\in\mathbb{T}_{n}$. Let $t^{\prime}\in\mathbb{T}_{n}$ be $k$-adjacent to $t$. We calculate $y_{il;t^{\prime}}^{\mathrm{c}}$ by the formula (3.8). When $i=k$, $y_{kl;t^{\prime}}^{\mathrm{c}}=s_{kl}^{-\sigma_{k;t}}(y_{k;t}^{\mathrm{g}})^{-1}\mathfrak{u}_{k;t}^{-1}$ also satisfies the same property. Assume $i\neq k$. Then, we have
(4.15) $y_{il;t^{\prime}}^{\mathrm{c}}$ $=s_{il}^{\sigma_{i;t}}y_{i;t}^{\mathrm{g}}(y_{k;t}^{\mathrm{g}})^{r_{k}[b_{ki;t}]_{+}}Z_{k;t}|_{\mathbb{S}_{0}}(y_{k;t}^{\mathrm{g}})^{-b_{ki;t}}$
$\qquad\times\mathfrak{u}_{i;t}\mathfrak{u}_{k;t}^{r_{k}[b_{ki;t}]_{+}}\left(\prod_{m=1}^{r_{k}}s_{km}\right)^{\sigma_{k;t}[b_{ki;t}]_{+}}\left(\frac{\prod_{m=1}^{r_{k}}(1\bar{\oplus}s_{km}^{\sigma_{k;t}}y_{k;t}^{\mathrm{g}}\mathfrak{u}_{k;t})}{Z_{k;t}|_{\mathbb{S}_{0}}(y_{k;t}^{\mathrm{g}})}\right)^{-b_{ki;t}}.$
Then, every factor on the second row of (4.15) belongs to $\mathbb{S}_{1}$ due to (4.5) and (4.13). Thus, the claim holds. ∎
Lemma 4.4.
For any $i=1,\dots,n$ and $t\in\mathbb{T}_{n}$, the product $\prod_{l=1}^{r_{i}}y_{il;t}^{\mathrm{c}}$ belongs to $\mathbb{S}$.
Proof.
This follows from (4.14). ∎
As a result, ${\bf\Upsilon}^{\mathrm{g}}$ is realized in $\mathbb{S}/{\sim}\cong\mathbb{P}$ by ${\bf\Upsilon}^{\mathrm{c}}$ as follows.
Theorem 4.5.
In $\mathbb{S}/{\sim}$, we have
(4.16) $y_{i;t}^{\mathrm{g}}]=\left[\prod_{l=1}^{r_{i}}y_{il;t}^{\mathrm{c}}\right]^{\frac{1}{r_{i}}},$
where the right hand side is the unique $r_{i}$th root of $\left[\prod_{l=1}^{r_{i}}y_{il;t}^{\mathrm{c}}\right]$ in $\mathbb{S}/{\sim}$.
Proof.
By (4.5) and (4.14), we have
(4.17) $\prod_{l=1}^{r_{i}}y_{il;t}^{\mathrm{c}}\sim(y_{i;t}^{\mathrm{g}})^{r_{i}}.$
Thus, $y_{i;t}^{\mathrm{g}}]$ is an $r_{i}$th root of $\prod_{l=1}^{r_{i}}y_{il;t}^{\mathrm{c}}]$. Also, by [FZ02], every semifield is torsion-free as a multiplicative abelian group. Therefore, it is unique. ∎
5. Realization of generalized cluster algebras
In this section, for a given GCA $\mathcal{A}^{\mathrm{g}}$, we construct a CA $\mathcal{A}$ whose subquotient $\mathcal{A}^{\mathrm{c}}/{\mathcal{I}}$ is isomorphic to $\mathcal{A}^{\mathrm{g}}$.
5.1. Realization of generalized cluster patterns
As in Section 4, let ${\bf\Sigma}^{\mathrm{g}}=\{({\bf x}^{\mathrm{g}}_{t},{\bf y}^{\mathrm{g}}_{t},{\bf Z}_{t},B_{t})\}_{t\in\mathbb{T}_{n}}$ be a generalized cluster pattern with the initial seed $\Sigma^{\mathrm{g}}_{t_{0}}$. Let $\overline{\mathbb{P}}=\mathbb{QP}_{\mathrm{sf}}({\bf s})$, and let $\overline{\mathcal{F}}$ be the field of the rational functions in $\mathcal{N}=r_{1}+\cdots+r_{n}$ variables with coefficients in $\mathbb{Q}\overline{\mathbb{P}}$. We take an ordinary seed $\Sigma_{t_{0}}^{\mathrm{c}}=({\bf x}_{t_{0}}^{\mathrm{c}},{\bf y}_{t_{0}}^{\mathrm{c}},\mathcal{B}_{t_{0}})$ of rank $\mathcal{N}$ with the ambient field $\overline{\mathcal{F}}$ such that its $Y$-seed $({\bf y}_{t_{0}}^{\mathrm{c}},\mathcal{B}_{t_{0}})$ is given by (4.11). Then, we consider the composite cluster pattern ${\bf\Sigma}^{\mathrm{c}}=\{({\bf x}_{t}^{\mathrm{c}},{\bf y}_{t}^{\mathrm{c}},\mathcal{B}_{t})\}_{t\in\mathbb{T}_{n}}$ with the initial seed $\Sigma_{t_{0}}^{\mathrm{c}}$, which is a subcollection of the ordinary cluster pattern ${\bf\Sigma}=\{({\bf x}_{t},{\bf y}_{t},\mathcal{B}_{t})\}_{t\in\mathbb{T}_{\mathcal{N}}}$ with the same initial seed $\Sigma_{t_{0}}=\Sigma_{t_{0}}^{\mathrm{c}}$.
For any $t\in\mathbb{T}_{n}$, we set
(5.1) $X_{i;t}=\prod_{l=1}^{r_{i}}x_{il;t}^{\mathrm{c}},\quad\hat{Y}_{i;t}=y_{i;t}^{\mathrm{g}}\prod_{j=1}^{r_{i}}X_{j;t}^{b_{ji;t}}.$
Then, for any $(i,l)\in\mathcal{D}$ and $t\in\mathbb{T}_{n}$, $\hat{y}-$variables $\hat{y}_{il;t}^{\mathrm{c}}$ of $\Sigma_{t}^{\mathrm{c}}$ is given by
[…]
We write ${\bf X}=(X_{1;t_{0}},\dots,X_{n;t_{0}})$, for simplicity.
In parallel to $\mathbb{S}$ in (4.6), we introduce the $\mathbb{ZS}$-subalgebra $\mathcal{S}\subset\overline{\mathcal{F}}$ as follows. For each $(i,l)\in\mathcal{D}$, let
(5.3) $\hat{e}_{il}=\sum_{1\leq m_{1}<m_{2}<\cdots<m_{l}\leq r_{i}}s_{im_{1}}s_{im_{2}}\cdots s_{im_{l}}\in\overline{\mathcal{F}}$
be the elementary symmetric polynomial in $s_{i1},\dots,s_{ir_{i}}$ of degree $l$ with respect to the addition ++ in $\overline{\mathcal{F}}$. Let $\hat{\bf e}=\{\hat{e}_{il}\mid(i,l)\in\mathcal{D}\}$. Let $\mathbb{QP}({\bf X})$ be the fraction field of $\mathbb{ZP}[{\bf X}]$. Then, the semifield homomorphism $\psi:\mathbb{S}\to\mathbb{P}$ in (4.7) can be extended to a $\mathbb{ZP}$-algebra homomorphism $\hat{\psi}:\mathbb{ZS}[\hat{\bf e},{\bf X}]\to\mathbb{ZP}[{\bf X}]$ by $X_{i;t_{0}}\mapsto X_{i;t_{0}}$ and $\hat{e}_{il}\mapsto\hat{z}_{il}$. Let
(5.4) $\mathcal{S}=\left\{\left.\frac{f}{g}\in\overline{\mathcal{F}}\ \right|\ f,g\in\mathbb{ZS}[\hat{\bf e},{\bf X}],\ \hat{\psi}(g)\neq 0\right\}.$
This is a $\mathbb{ZP}$-subalgebra of $\overline{\mathcal{F}}$. One can extend the above $\hat{\psi}$ to a $\mathbb{ZP}$-algebra homomorphism $\hat{\psi}:\mathcal{S}\to\mathbb{QP}({\bf X})$ by $\psi(f/g)=\psi(f)/\psi(g)$. Then, the following fact holds.
Lemma 5.1.
The quotient $\mathcal{S}/{\mathrm{Ker}\,\hat{\psi}}$ is a field. Moreover, the $\mathbb{ZP}$-algebra homomorphism $\mathbb{QP}({\bf X})\to\mathcal{S}/{\mathrm{Ker}\,\hat{\psi}}$ defined by $f\mapsto[f]$ is a field isomorphism.
Proof.
The composition $\mathbb{QP}({\bf X})\hookrightarrow\mathcal{S}\overset{\hat{\psi}}{\to}\mathbb{QP}({\bf X})$ is the identity map, where $\mathbb{QP}({\bf X})\hookrightarrow\mathcal{S}$ is the canonical inclusion. Thus, $\hat{\psi}$ is surjective. Then, by the fundamental theorem on homomorphisms, the claim holds. ∎
Let $\equiv$ denote the equivalence in $\mathcal{S}$ modulo the ideal $\mathrm{Ker}\,\hat{\psi}$. Now, we have an analogous relation (4.3) in $\mathcal{S}$.
Lemma 5.2 (cf. [RW25, Lem. 29]).
For any $u\in\mathcal{S}$ and $k=1,\dots,n$, the products $\prod_{l=1}^{r_{k}}(1+s_{kl}u)$ and $\prod_{l=1}^{r_{k}}(1+s_{kl}^{-1}u)$ belong to $\mathcal{S}$. Moreover, we have
(5.5) $\prod_{l=1}^{r_{k}}(1+s_{kl}^{\sigma_{k;t}}u)\equiv Z_{k;t}(u).$
Proof.
They are proved in the same way as Lemma 4.2 and (4.13). ∎
Lemma 5.3.
Every $X_{i;t}$ belongs to $\mathcal{S}$, and it holds that $\hat{\psi}(X_{i;t})\neq 0$.
Proof.
We prove the claim by the induction on $t\in\mathbb{T}_{n}$. For $t=t_{0}$, the claim holds by definition. Suppose that the claim holds for some $t\in\mathbb{T}_{n}$, and let $t^{\prime}\in\mathbb{T}_{n}$ be $k$-adjacent to $t$. By (3.7), (4.14), and (5.2), we have
(5.6) $X_{k;t^{\prime}}=\frac{1}{X_{k;t}}\left(\prod_{j=1}^{n}X_{j;t}^{[-b_{jk;t}]_{+}}\right)^{r_{k}}\frac{\prod_{l=1}^{r_{k}}(1+s_{kl}^{\sigma_{k;t}}\hat{Y}_{k;t}\mathfrak{u}_{k;t})}{\prod_{l=1}^{r_{k}}(1\bar{\oplus}s_{kl}^{\sigma_{k;t}}y_{k;t}^{\mathrm{g}}\mathfrak{u}_{k;t})}.$
By the assumption, the first and the second factors on the right hand side satisfy the claim. Also, by Lemmas 4.2 and 5.2, the third factor belongs to $\mathcal{S}$, and the image of it by $\hat{\psi}$ is nonzero. ∎
Thanks to Lemma 5.1, we have the field isomorphism $\Phi:\mathcal{F}=\mathbb{QP}({\bf x}_{t_{0}}^{\mathrm{g}})\to\mathcal{S}/{\mathrm{Ker}\,}\hat{\psi}$ by
(5.7) $\Phi(x_{i;t_{0}}^{\mathrm{g}})=[X_{i;t_{0}}],\quad\Phi(p)=[p]\quad(p\in\mathbb{P}).$
Now, the generalized cluster pattern ${\bf\Sigma}^{\mathrm{g}}$ is realized by the composite cluster pattern ${\bf\Sigma}^{\mathrm{c}}$ as follows.
Theorem 5.4 (cf. [RW25, Thm. 30]).
By the above isomorphism $\Phi:\mathcal{F}\to\mathcal{S}/\mathrm{Ker}\,\hat{\psi}$, we have
(5.8) $y_{i;t}^{\mathrm{g}}\mapsto\left[\prod_{l=1}^{r_{i}}y_{il;t}^{\mathrm{c}}\right]^{\frac{1}{r_{i}}},\quad x_{i;t}^{\mathrm{g}}\mapsto[X_{i;t}].$
Proof.
For the $y$-variables, we have already shown the claim in Theorem 4.5. For the $x$-variables, if the cliam holds for some $t\in\mathbb{T}_{n}$, by (5.1), we have $\Phi(\hat{y}_{i;t}^{\mathrm{g}})=[\hat{Y}_{i;t}]$. Then, by applying (4.13) and (5.5) to (5.6), we obtain the claim by the induction on $t\in\mathbb{T}_{n}$. ∎
5.2. Realization of GCAs
Let $\mathcal{A}^{\mathrm{g}}=\mathcal{A}({\bf\Sigma}^{\mathrm{g}})$ be the GCA of ${\bf\Sigma}^{\mathrm{g}}$, and let $\mathcal{A}=\mathcal{A}({\bf\Sigma})$ be the CA of $\bf\Sigma$}. Consider the following $\mathbb{ZS}$-subalgebra of $\mathcal{A}$:
(5.9) $\mathcal{A}^{\mathrm{c}}=\mathbb{ZS}[X_{i;t}\mid i=1,\dots,n,\ t\in\mathbb{T}_{n}]\subset\mathcal{A}.$
By Lemma 5.3, we also have $\mathcal{A}^{\mathrm{c}}\subset\mathcal{S}$. Let $\mathcal{I}=\mathrm{Ker}\,\hat{\psi}\cap\mathcal{A}^{\mathrm{c}}$, and consider its quotient $\mathcal{A}^{\mathrm{c}}/{\mathcal{I}}$.
We set $\tilde{\mathbb{S}}=\mathbb{S}/{\sim}$. All relevant algebras and their relations are summarized in the following diagram.
(5.10) $\overline{\mathcal{F}}=\mathbb{Q}\overline{\mathbb{P}}(\mathbf{x}_{t_{0}})$$\mathcal{S}$$\mathcal{A}=\mathbb{Z}\overline{\mathbb{P}}[x_{i;t}|i,t]$$\mathcal{F}=\mathbb{Q}\mathbb{P}(\mathbf{x}_{t_{0}}^{\mathrm{g}})$$\mathcal{S}/\mathrm{Ker}\,\hat{\psi}$$\mathcal{A}^{\mathrm{c}}+\mathrm{Ker}\,\hat{\psi}$$\mathcal{A}^{\mathrm{c}}=\mathbb{Z}\mathbb{S}[X_{i;t}|i,t]$$\mathcal{A}^{\mathrm{g}}=\mathbb{Z}\mathbb{P}[x_{i;t}^{\mathrm{g}}|i,t]$$\mathbb{Z}\tilde{\mathbb{S}}[[X_{i;t}]|i,t]$$\mathcal{A}^{\mathrm{c}}/\mathcal{I}\simeq\simeq\simeq\subset\subset\subset\subset\subset\subset\supset$
The following is the main result, which is parallel to the result of [RW25].
Theorem 5.5 (cf. [RW25, Thm. 1]).
The $\mathbb{ZP}$-algebra homomorphism $\Psi:\mathcal{A}^{\mathrm{g}}\to\mathcal{A}^{\mathrm{c}}/\mathcal{I}$ defined by $x_{i;t}^{\mathrm{g}}\mapsto[X_{i;t}]$ and $p\mapsto[p]$ ($p\in\mathbb{P}$) is an isomorphism. Therefore, $\mathcal{A}^{\mathrm{g}}$ is isomorphic to a subquotient $\mathcal{A}^{\mathrm{c}}/\mathcal{I}$ of $\mathcal{A}$.
Proof.
In the bottom line of (5.10), the first isomorphism $\mathcal{A}\cong\mathbb{Z}\tilde{\mathbb{S}}[[X_{i;t}]\mid i,t]$, $x_{i;t}^{\mathrm{g}}\mapsto[X_{i;t}]$ is obtained by Theorem 5.4. The second isomorphism $\mathbb{Z}\tilde{\mathbb{S}}[[X_{i;t}]\mid i,t]\cong\mathcal{A}^{\mathrm{c}}/(\mathcal{I})$, $X_{i;t}]\mapsto[X_{i;t}]$ is due to the standard isomorphism theorem. The map $\Psi:\mathcal{A}^{\mathrm{g}}\to\mathcal{A}^{\mathrm{c}}/\mathcal{I}$ is the composition of them. ∎
6. Relations for $C$-matrices, $G$-matrices, and $F$-polynomials
In this section, we present the relations between the $C$-matrices, the $G$-matrices, and the $F$-polynomials of a generalized cluster pattern $\mathbf{\Sigma}^{\mathrm{g}}$ and those of the associated composite cluster pattern $\mathbf{\Sigma}^{\mathrm{c}}$. These relations are compatible with the correspondence (5.8).
6.1. Separation formulas
In parallel to the ordinary case [FZ07], to any generalized cluster pattern $\mathbf{\Sigma}^{\mathrm{g}}$ with $y$-variables, one can associate the collections of the $C$-matrices $\{C^{\mathrm{g}}_{t}\}_{t\in\mathbb{T}_{n}}$, the $G$-matrices $\{G^{\mathrm{g}}_{t}\}_{t\in\mathbb{T}_{n}}$, and the $F$-polynomials $\{F^{\mathrm{g}}_{i;t}\}_{i=1,\dots,n;t\in\mathbb{T}_{n}}$ [NAK15b]. Here, each $F$-polynomial $F^{\mathrm{g}}_{i;t}(\mathbf{y},\mathbf{z})$ is a polynomial in formal variables $\mathbf{y}=(y_{i})_{i=1}^{n}$ and $\mathbf{z}=(z_{il})_{(i,l)\in\mathcal{D}^{\prime}}$ with (nonnegative) integer coefficients, where $\mathcal{D}^{\prime}=\{(i,l)\mid i=1,\dots,n;\ l=1,\dots,r_{i}-1\}$. Then, the $x$- and $y$-variables of $\mathbf{\Sigma}^{\mathrm{g}}$ are expressed by the following separation formulas [NAK15b, NR16]:
(6.1) $x^{\mathrm{g}}_{i;t}$ $=\left(\prod_{j=1}^{n}(x^{\mathrm{g}}_{j})^{g^{\mathrm{g}}_{ji;t}}\right)\frac{F^{\mathrm{g}}_{i;t}(\hat{\mathbf{y}}^{\mathrm{g}},\hat{\mathbf{z}})}{F^{\mathrm{g}}_{i;t}|_{\mathbb{P}_{0}}(\mathbf{y}^{\mathrm{g}},\mathbf{z})},$
(6.2) $y^{\mathrm{g}}_{i;t}$ $=\prod_{j=1}^{n}(y^{\mathrm{g}}_{j})^{c^{\mathrm{g}}_{ji;t}}\prod_{j=1}^{n}F^{\mathrm{g}}_{j;t}|_{\mathbb{P}_{0}}(\mathbf{y}^{\mathrm{g}},\mathbf{z})^{b_{ji;t}},$
where ${\bf x}^{\mathrm{g}}$, ${\bf y}^{\mathrm{g}}$, and $\hat{\bf y}^{\mathrm{g}}$ are the initial $x$-, $y$-, $\hat{y}-$variables of $\mathbf{\Sigma}^{\mathrm{g}}$, and $\hat{\mathbf{z}}$ and $\mathbf{z}$} are the coefficients of the initial exchange polynomials $Z_{i}(u)$ and their images of $\pi_{\mathbb{P}_{0}}$. There is an unfortunate coincidence of notations $g$ and $\mathrm{g}$, and also, $c$ and $\mathrm{c}$ for entirely different meanings, and we ask the readers to carefully distinguish them.
[…]
(6.3) $x^{\mathrm{c}}_{il;t}$ $=\biggl(\prod_{(j,m)\in\mathcal{D}}(x^{\mathrm{c}}_{jm})^{g^{\mathrm{c}}_{jm,il;t}}\biggr)\frac{F^{\mathrm{c}}_{il;t}(\hat{\mathbf{y}}^{\mathrm{c}})}{F^{\mathrm{c}}_{il;t}|_{\overline{\mathbb{P}}}(\mathbf{y}^{\mathrm{c}})},$
(6.4) $y^{\mathrm{c}}_{il;t}$ $=\prod_{(j,m)\in\mathcal{D}}(y^{\mathrm{c}}_{jm})^{c^{\mathrm{c}}_{jm,il;t}}\prod_{(j,m)\in\mathcal{D}}F^{\mathrm{c}}_{jm;t}|_{\overline{\mathbb{P}}}(\mathbf{y}^{\mathrm{c}})^{b_{ji;t}},$
where $\mathbf{x}^{\mathrm{c}}$, $\mathbf{y}^{\mathrm{c}}$, and $\hat{\mathbf{y}}^{\mathrm{c}}$ are the initial $x$-, $y$-, $\hat{y}-$variables of $\mathbf{\Sigma}^{\mathrm{c}}$.
6.2. Relations for $C$-matrices and $G$-matrices
For any $i,j=1,\dots,n$, set $\delta_{ij}=1$ if $i=j$ and $\delta_{ij}=0$ if $i\neq j$. The $C$-matrices and the $G$-matrices are determined by the initial condition $c_{ij;t_{0}}^{\mathrm{g}}=g_{ij;t_{0}}^{\mathrm{g}}=\delta_{ij}$ and the following recursions for each pair of $k$-adjacent vertices $t,t^{\prime}\in\mathbb{T}_{n}$ [NAK15b, NR16]:
(6.5) $c_{ij;t^{\prime}}^{\mathrm{g}}$ $=\begin{cases}-c_{ik;t}^{\mathrm{g}}&j=k,\\ c_{ij;t}^{\mathrm{g}}+r_{k}(c_{ik;t}^{\mathrm{g}}[b_{kj;t}]_{+}+[-c_{ik;t}^{\mathrm{g}}]_{+}b_{kj;t})&j\neq k,\end{cases}$
(6.6) $g_{ij;t^{\prime}}^{\mathrm{g}}$ $=\begin{cases}\displaystyle{-g_{ik;t}^{\mathrm{g}}+r_{k}\sum_{a=1}^{n}\left(g_{ia;t}^{\mathrm{g}}[-b_{ak;t}]_{+}-b_{ia;t_{0}}[-c_{ak;t}^{\mathrm{g}}]_{+}\right)}&j=k,\\ g_{ij;t}^{\mathrm{g}}&j\neq k.\end{cases}$
For any $(i,l),(j,m)\in\mathcal{D}$, let $\delta_{il,jm}=1$ if $(i,l)=(j,m)$ and $\delta_{il,jm}=0$ if $(i,l)\neq(j,m)$. Recall that the sign $\sigma_{i;t}$ is the one in (4.12). Then, the following fact holds.
Lemma 6.1.
Let $\tilde{c}_{il,jm;t}^{\mathrm{c}}=c_{il,jm;t}^{\mathrm{c}}-\sigma_{j;t}\delta_{il,jm}$ and $\tilde{g}_{il,jm;t}^{\mathrm{c}}=g_{il,jm;t}^{\mathrm{c}}-\sigma_{j;t}\delta_{il,jm}$. Then, they are independent of $l=1,\dots,r_{i}$ and $m=1,\dots,r_{j}$.
Proof.
For $t=t_{0}$, since $\tilde{c}_{il,jm;t_{0}}^{\mathrm{c}}=\tilde{g}_{il,jm;t_{0}}^{\mathrm{c}}=0$, the claim holds. Suppose that the claim holds for some $t\in\mathbb{T}_{n}$. Let $t^{\prime}\in\mathbb{T}_{n}$ be $k$-adjacent to $t$. Consider the $C$-matrix $C_{t^{\prime}}^{\mathrm{c}}$. By applying (6.5) with ${\bf r}=(1,\dots,1)$ repeatedly, we have
(6.7) $c_{il,jm;t^{\prime}}^{\mathrm{c}}=\begin{cases}-c_{il,km;t}^{\mathrm{c}}&j=k,\\ \displaystyle{c_{il,jm;t}^{\mathrm{c}}+\sum_{p=1}^{r_{k}}(c_{il,kp;t}^{\mathrm{c}}[b_{kj;t}]_{+}+[-c_{il,kp;t}^{\mathrm{c}}]_{+}b_{kj;t})}&j\neq k.\end{cases}$
When $j=k$, the claim holds because $\tilde{c}_{il,km;t^{\prime}}^{\mathrm{c}}=-\tilde{c}_{il,km;t}^{\mathrm{c}}$. Consider the case $j\neq k$. By the assumption, the sums $\sum_{p=1}^{r_{k}}c_{il,kp;t}^{\mathrm{c}}$ and $\sum_{p=1}^{r_{k}}[-c_{il,kp;t}^{\mathrm{c}}]_{+}$ are independent of $l=1,\dots,n$. Therefore, we have the claim by (6.7). For the $G$-matrix $G_{t^{\prime}}^{\mathrm{c}}$, when $j\neq k$, we have $\tilde{g}_{il,jm;t^{\prime}}^{\mathrm{c}}=\tilde{g}_{il,jm;t}^{\mathrm{c}}$. Thus, the claim holds. When $j=k$, we have
(6.8) $g_{il,km;t^{\prime}}^{\mathrm{c}}=-g_{il,km;t}^{\mathrm{c}}+\sum_{a=1}^{n}\sum_{p=1}^{r_{a}}\left(g_{il,ap;t}^{\mathrm{c}}[-b_{ak;t}]_{+}-b_{ia;t_{0}}[-c_{ap,km;t}^{\mathrm{c}}]_{+}\right).$
Then, the rest of the argument is similar to the above. ∎
In view of Lemma 6.1, we set
(6.9) $\tilde{c}_{ij;t}^{\mathrm{c}}=c_{il,jm;t}^{\mathrm{c}}-\sigma_{j;t}\delta_{il,jm},\quad\tilde{g}_{ij;t}^{\mathrm{c}}=g_{il,jm;t}^{\mathrm{c}}-\sigma_{j;t}\delta_{il,jm},$
which are independent of $l$ and $m$. Then, we have the following relations.
Proposition 6.2.
The following equalities hold:
(6.10) $c_{ij;t}^{\mathrm{g}}$ $=\sum_{l=1}^{r_{i}}c_{il,jm_{0};t}^{\mathrm{c}}=\frac{r_{i}}{r_{j}}\sum_{m=1}^{r_{j}}c_{il_{0},jm;t}^{\mathrm{c}}=r_{i}\tilde{c}_{ij;t}^{\mathrm{c}}+\sigma_{j;t}\delta_{ij},$
(6.11) $g_{ij;t}^{\mathrm{g}}$ $=\sum_{m=1}^{r_{j}}g_{il_{0},jm;t}^{\mathrm{c}}=\frac{r_{j}}{r_{i}}\sum_{l=1}^{r_{i}}g_{il,jm_{0};t}^{\mathrm{c}}=r_{j}\tilde{g}_{ij;t}^{\mathrm{c}}+\sigma_{j;t}\delta_{ij},$
where $l_{0}=1,\dots,r_{i}$ and $m_{0}=1,\dots,r_{j}$ are arbitrary.
Proof.
For $t=t_{0}$, the claim holds by definition. Suppose that the claim holds for some $t\in\mathbb{T}_{n}$, and let $t^{\prime}\in\mathbb{T}_{n}$ be $k$-adjacent to $t$. Consider the $C$-matrix $C_{t^{\prime}}^{\mathrm{c}}$. By (6.7), the claim holds when $j=k$. When $j\neq k$, we have
(6.12) $\sum_{l=1}^{r_{i}}c_{il,jm;t^{\prime}}^{\mathrm{c}}=\sum_{l=1}^{r_{i}}c_{il,jm;t}^{\mathrm{c}}+\sum_{p=1}^{r_{k}}\left(\sum_{l=1}^{r_{i}}c_{il,kp;t}^{\mathrm{c}}[b_{kj;t}]_{+}+\sum_{l=1}^{r_{i}}[-c_{il,kp;t}^{\mathrm{c}}]_{+}b_{kj;t}\right).$
Since $\sigma_{j;t}\in\{\pm 1\}$, $c_{il,kp;t}^{\mathrm{c}}$ ($l=1,\dots,r_{i}$) are either all nonnegative or all nonpositive. Thus, we have
(6.13) $\sum_{l=1}^{r_{i}}[-c_{il,kp;t}^{\mathrm{c}}]_{+}=[-c_{ik;t}^{\mathrm{g}}]_{+}.$
Then, we have
[…]
Thus, the first equality in (6.10) holds. Then, by (6.9), we also obtain the others in (6.10). For the $G$-matrix $G_{t^{\prime}}^{\mathrm{c}}$, when $j\neq k$, the claim holds because $g_{il,jm;t^{\prime}}^{\mathrm{c}}=g_{il,jm;t}^{\mathrm{c}}$. When $j=k$, by (6.8) and (6.13), we have
[…]
Thus, the first equality in (6.11) holds. Then, by (6.9), we also obtain the others in (6.11). ∎
6.3. Relation for $F$-polynomials
For each $i=1,\dots,n$, let
(6.16) $\mathcal{Z}_{i;t_{0}}(u)=1+z_{i1}u+z_{i2}u^{2}+\cdots+z_{ir_{i}-1}u^{r_{i}-1}+u^{r_{i}},$
and we recursively define $\mathcal{Z}_{i;t}(u)$ ($t\in\mathbb{T}_{n}$) by $\mathcal{Z}_{i;t^{\prime}}(u)=\mathcal{Z}_{i;t}(u)$ if $i\neq k$ and $\mathcal{Z}_{k;t^{\prime}}(u)=\mathcal{Z}_{k;t}^{\mathrm{rec}}(u)$ for any pair of $k$-adjacent vertices $t,t^{\prime}\in\mathbb{T}_{n}$. The $F$-polynomials are determined by the initial condition $F_{i;t_{0}}=1$ ($i=1,\dots,n$) and the following recursion for each pair of $k$-adjacent vertices $t,t^{\prime}\in\mathbb{T}_{n}$ [NAK15b, NR16]:
(6.17) $F_{i;t^{\prime}}^{\mathrm{g}}({\bf y},{\bf z})=\begin{cases}F_{k;t}^{\mathrm{g}}({\bf y},{\bf z})^{-1}\cdot M({\bf y},{\bf z})i=k,\\ F_{i;t}^{\mathrm{g}}({\bf y},{\bf z})&i\neq k,\end{cases}$
where
(6.18) $M({\bf y},{\bf z})=\left(\prod_{j=1}^{n}y_{j}^{[-c_{jk;t}^{\mathrm{g}}]_{+}}F_{j;t}^{\mathrm{g}}({\bf y},{\bf z})^{[-b_{jk;t}]_{+}}\right)^{r_{k}}\mathcal{Z}_{k;t}\left(\prod_{j=1}^{n}y_{j}^{c_{jk;t}^{\mathrm{g}}}F_{j;t}^{\mathrm{g}}({\bf y},{\bf z})^{b_{jk;t}}\right).$
Recall that each $F$-polynomial of the composite cluster pattern ${\bf\Sigma}^{\mathrm{c}}$ is a polynomial in $\tilde{\bf y}=(\tilde{y}_{il})_{(i,l)\in\mathcal{D}}$. Let $\mathfrak{S}(\mathcal{D})$ be the symmetric group of $\mathcal{D}$. For any $\sigma\in\mathfrak{S}(\mathcal{D})$, we write
(6.19) $\tilde{\bf y}^{\sigma}=(\tilde{y}_{\sigma(i,l)})_{(i,l)\in\mathcal{D}}.$
For each $j=1,\dots,n$, let $\mathcal{D}_{j}=\{(j,m)\in\mathcal{D}\mid m=1,\dots,r_{j}\}$. Then, the following fact holds.
Lemma 6.3.
Let $(i,l)\in\mathcal{D}$ and $t\in\mathbb{T}_{n}$. Let $\sigma\in\mathfrak{S}(\mathcal{D})$ satisfy $\sigma(\mathcal{D}_{j})=\mathcal{D}_{j}$ for any $j=1,\dots,n$. Then, we have
[…]
Proof.
Suppose that the claim holds for some $t\in\mathbb{T}_{n}$. Let $t^{\prime}\in\mathbb{T}_{n}$ be $k$-adjacent to $t$. Set
[…]
We introduce formal variables $s_{jm}$ ($(j,m)\in\mathcal{D}$), and we set
(6.22) ${\bf s}=(s_{jm})_{(j,m)\in\mathcal{D}},\quad{\bf s}{\bf y}=(s_{jm}y_{j})_{(j,m)\in\mathcal{D}}.$
Then, by Lemma 6.3, for each $i$ and $j$, the product $\prod_{l=1}^{r_{i}}F_{il;t}^{\mathrm{c}}({\bf sy})$ is symmetric with respect to $s_{j1},\dots,s_{jr_{j}}$. In particular, it can be expressed as a polynomial in $\bf y$} and ${\bf e}=(e_{jm})_{(j,m)\in\mathcal{D}}$, where $e_{jm}$ is the elementary symmetric polynomial in $s_{j1},\dots,s_{jr_{j}}$ of degree $m$. In this expression, we apply the substitution $e_{jm}=z_{jm}$ ($(j,m)\in\mathcal{D}^{\prime}$) and $e_{jr_{j}}=1$ ($j=1,\dots,n$). The resulting polynomial in $\bf y$} and $\bf z$} is denoted by $\prod_{l=1}^{r_{i}}F_{il;t}^{\mathrm{c}}({\bf sy})|_{{\bf e}={\bf z}}$, for simplicity, where we slightly abuse the symbol $\bf z$}.
Let $\tau_{il,im}\in\mathfrak{S}(\mathcal{D})$ be the transposition of $(i,l)$ and $(i,m)$ for $l\neq m$ and the identity for $l=m$. Then, we have the following relations.
Proposition 6.4.
For arbitrary $l_{0}=1,\dots,r_{i}$, the following equalities hold:
(6.23) $F_{i;t}^{\mathrm{g}}({\bf y},{\bf z})=\left.\prod_{l=1}^{r_{i}}F_{il;t}^{\mathrm{c}}({\bf s}{\bf y})\right|_{{\bf e}={\bf z}}=\left.\prod_{m=1}^{r_{i}}F_{il_{0};t}^{\mathrm{c}}\left(({\bf s}{\bf y})^{\tau_{il_{0},im}}\right)\right|_{{\bf e}={\bf z}}.$
Proof.
The second equality follows from Lemma 6.3. We show the first one. For $t=t_{0}$, the claim holds by definition. Suppose that the claim holds for some $t\in\mathbb{T}_{n}$, and let $t^{\prime}\in\mathbb{T}_{n}$ be $k$-adjacent to $t$. Then, we have $F_{k;t^{\prime}}^{\mathrm{g}}({\bf y},{\bf z})=F_{k;t}^{\mathrm{g}}({\bf y},{\bf z})^{-1}M({\bf y},{\bf z})$, where $M$ is the polynomial defined in (6.18). On the other hand, we have $\prod_{l=1}^{r_{k}}F_{kl;t^{\prime}}^{\mathrm{c}}({\bf s}{\bf y})=(\prod_{l=1}^{r_{k}}F_{kl;t}^{\mathrm{c}}({\bf s}{\bf y}))^{-1}P({\bf s}{\bf y})Q({\bf s}{\bf y})$, where
(6.24) $P({\bf sy})$ $=\prod_{l=1}^{r_{k}}\prod_{(j,m)\in\mathcal{D}}(s_{jm}y_{j})^{[-c_{jm,kl;t}^{\mathrm{c}}]_{+}}F_{jm;t}^{\mathrm{c}}({\bf sy})^{[-b_{jk;t}]_{+}},$
(6.25) $Q({\bf sy})$ $=\prod_{l=1}^{r_{k}}\biggl(1+\prod_{(j,m)\in\mathcal{D}}(s_{jm}y_{j})^{c_{jm,kl;t}^{\mathrm{c}}}F_{jm;t}^{\mathrm{c}}({\bf sy})^{b_{jk;t}}\biggr).$
Comparing the first and the third terms in (6.10), we have
(6.26) $\prod_{l=1}^{r_{k}}\prod_{(j,m)\in\mathcal{D}}s_{jm}^{[-c_{jm,kl;t}^{\mathrm{c}}]_{+}}=\prod_{(j,m)\in\mathcal{D}}s_{jm}^{\frac{r_{k}}{r_{j}}[-c_{jk;t}^{\mathrm{g}}]_{+}}=\prod_{j=1}^{n}e_{jr_{j}}^{\frac{r_{k}}{r_{j}}[-c_{jk;t}^{\mathrm{c}}]_{+}}.$
Thus, by (6.10), we obtain
(6.27) $P({\bf s}{\bf y})|_{{\bf e}={\bf z}}=\prod_{j=1}^{n}y_{j}^{r_{k}[-c_{jk;t}^{\mathrm{g}}]_{+}}F_{j;t}^{\mathrm{g}}({\bf y},{\bf z})^{r_{k}[-b_{jk;t}]_{+}}.$
Meanwhile, by (6.9), we have
(6.28) $\prod_{(j,m)\in\mathcal{D}}s_{jm}^{c_{jm,kl;t}^{\mathrm{c}}}=s_{kl}^{\sigma_{k;t}}\prod_{j=1}^{n}e_{jr_{j}}^{\tilde{c}_{jk;t}^{\mathrm{c}}}.$
Thus, we have
[…]
$=\mathcal{Z}_{k;t}\left(\prod_{j=1}^{n}y_{j}^{c_{jk;t}^{\mathrm{g}}}F_{j;t}^{\mathrm{g}}({\bf y},{\bf z})^{b_{jk;t}}\right),$
where the second equality is shown by the same way as (4.13). By (6.27) and (6.29), we have $P({\bf s}{\bf y})Q({\bf s}{\bf y})|_{{\bf e}={\bf z}}=M({\bf y},{\bf z})$. Thus, the claim holds. ∎
Remark 6.5.
The equality (6.23) does not directly imply the positivity of the coefficients in $F_{i;t}^{\mathrm{g}}$ (the Laurent positivity of $x_{i;t}^{\mathrm{g}}$ proved by [BLM25a]) from the one of $F_{il;t}^{\mathrm{c}}$ due to possible signs such as $s_{1}^{2}+s_{2}^{2}=(s_{1}+s_{2})^{2}-2s_{1}s_{2}$.
6.4. Examples
Let us give some examples of Propositions 6.2 and 6.4. We consider generalized cluster patterns of rank $2$. Let $t_{0}$ be the initial vertex, and consider the following vertices:
(6.30) $t_{0}$$t_{1}$$t_{2}$$t_{3}$112211
(1) We consider the case
(6.31) $\begin{gathered}B_{t_{0}}=\left(\begin{matrix}0&-1\\ 1&0\end{matrix}\right),\ Z_{1;t_{0}}(u)=1+\hat{z}u+u^{2},\ Z_{2;t_{0}}(u)=1+u,\\ \mathcal{B}_{t_{0}}=\left(\begin{array}[]{cc|c}0&0&-1\\ 0&0&-1\\ \hline\cr 1&1&0\end{array}\right).\end{gathered}$
The $C$-matrices, the $G$-matrices, and the $F$-polynomials are given in Table 1.
$t$ $C^{\mathrm{g}}_{t}$ $C^{\mathrm{c}}_{t}$ $G_{t}^{\mathrm{g}}$ $G_{t}^{\mathrm{c}}$
$t_{0}$ $\tiny\left(\begin{array}[]{cc}1&0\\ 0&1\end{array}\right)$} $\tiny\left(\begin{array}[]{cc|c}1&0&0\\ 0&1&0\\ \hline\cr 0&0&1\end{array}\right)$} $\tiny\left(\begin{array}[]{cc}1&0\\ 0&1\end{array}\right)$} $\tiny\left(\begin{array}[]{cc|c}1&0&0\\ 0&1&0\\ \hline\cr 0&0&1\end{array}\right)$}
$t_{1}$ $\tiny\left(\begin{array}[]{cc}-1&0\\ 0&1\end{array}\right)$} $\tiny\left(\begin{array}[]{cc|c}-1&0&0\\ 0&-1&0\\ \hline\cr 0&0&1\end{array}\right)$} $\tiny\left(\begin{array}[]{cc}-1&0\\ 0&1\end{array}\right)$} $\tiny\left(\begin{array}[]{cc|c}-1&0&0\\ 0&-1&0\\ \hline\cr 0&0&1\end{array}\right)$}
$t_{2}$ $\tiny\left(\begin{array}[]{cc}-1&0\\ 0&-1\end{array}\right)$} $\tiny\left(\begin{array}[]{cc|c}-1&0&0\\ 0&-1&0\\ \hline\cr 0&0&-1\end{array}\right)$} $\tiny\left(\begin{array}[]{cc}-1&0\\ 0&-1\end{array}\right)$} $\tiny\left(\begin{array}[]{cc|c}-1&0&0\\ 0&-1&0\\ \hline\cr 0&0&-1\end{array}\right)$}
$t_{3}$ $\tiny\left(\begin{array}[]{cc}-1&-2\\ 0&-1\end{array}\right)$} $\tiny\left(\begin{array}[]{cc|c}1&0&-1\\ 0&1&-1\\ \hline\cr 0&0&-1\end{array}\right)$} $\tiny\left(\begin{array}[]{cc}1&0\\ -2&-1\end{array}\right)$} $\tiny\left(\begin{array}[]{cc|c}1&0&0\\ 0&1&0\\ \hline\cr-1&-1&-1\end{array}\right)$}
$t$ $F^{\mathrm{g}}_{i;t}$ $F^{\mathrm{c}}_{il;t}$
$t_{0}$ $\begin{array}[]{c}1\\ 1\end{array}$ $\begin{array}[]{c}1\\ 1\\ 1\end{array}$
$t_{1}$ $\begin{array}[]{c}1+zy_{1}+y_{1}^{2}\\ 1\end{array}$ $\begin{array}[]{c}1+y_{11}\\ 1+y_{12}\\ 1\end{array}$
$t_{2}$ $\begin{array}[]{c}1+zy_{1}+y_{1}^{2}\\ 1+y_{2}+zy_{1}y_{2}+y_{1}^{2}y_{2}\end{array}$ $\begin{array}[]{c}1+y_{11}\\ 1+y_{12}\\ 1+y_{21}+(y_{11}+y_{12})y_{21}+y_{11}y_{12}y_{21}\end{array}$
$t_{3}$ $\begin{array}[]{c}1+2y_{2}+y_{2}^{2}+zy_{1}y_{2}+zy_{1}y_{2}^{2}+y_{1}^{2}y_{2}^{2}\\ 1+y_{2}+zy_{1}y_{2}+y_{1}^{2}y_{2}\end{array}$ $\begin{array}[]{c}1+y_{21}+y_{12}y_{21}\\ 1+y_{21}+y_{11}y_{21}\\ 1+y_{21}+(y_{11}+y_{12})y_{21}+y_{11}y_{12}y_{21}\end{array}$
Table 1. $C$- and $G$-matrices and $F$-polynomials for Case (1). In each box on the column of $F_{i;t}^{\mathrm{g}}$, we put $F_{1;t}^{\mathrm{g}}$ and $F_{2;t}^{\mathrm{g}}$, and, on the column of $F_{il;t}^{\mathrm{c}}$, we put $F_{11;t}^{\mathrm{c}}$, $F_{12;t}^{\mathrm{c}}$, and $F_{21;t}^{\mathrm{c}}$ in this order.
The equality (6.23) holds for $i=1$ and $t=t_{3}$ as follows:
(6.32) $\ F_{11;t_{3}}^{\mathrm{c}}(s_{11}y_{1},s_{12}y_{1},s_{21}y_{2})F_{12;t_{3}}^{\mathrm{c}}(s_{11}y_{1},s_{12}y_{1},s_{21}y_{2})|_{{\bf e}={\bf z}}$
$=$ $\ \{1+2s_{21}y_{2}+s_{21}^{2}y_{2}^{2}+(s_{11}+s_{12})s_{21}y_{1}y_{2}$
$\qquad+(s_{11}+s_{12})s_{21}^{2}y_{1}y_{2}^{2}+s_{11}s_{12}s_{21}^{2}y_{1}^{2}y_{2}^{2}\}|_{{\bf e}={\bf z}}$
$=$ $\ F_{1;t_{3}}^{\mathrm{g}}(y_{1},y_{2},z).$
(2) To illustrate Proposition 6.2 more evidently, we consider the case where the matrix $\mathcal{B}_{t_{0}}$ is the one in (3.2). The $C$-matrices and the $G$-matrices are given in Table 2.
$t$ $C_{t}^{\mathrm{g}}$ $C_{t}^{\mathrm{c}}$ $G_{t}^{\mathrm{g}}$ $G_{t}^{\mathrm{c}}$
$t_{0}$ $\tiny\left(\begin{array}[]{cc}1&0\\ 0&1\end{array}\right)$} $\tiny\left(\begin{array}[]{cc|ccc}1&0&0&0&0\\ 0&1&0&0&0\\ \hline\cr 0&0&1&0&0\\ 0&0&0&1&0\\ 0&0&0&0&1\end{array}\right)$} $\tiny\left(\begin{array}[]{cc}1&0\\ 0&1\end{array}\right)$} $\tiny\left(\begin{array}[]{cc|ccc}1&0&0&0&0\\ 0&1&0&0&0\\ \hline\cr 0&0&1&0&0\\ 0&0&0&1&0\\ 0&0&0&0&1\end{array}\right)$}
$t_{1}$ $\tiny\left(\begin{array}[]{cc}-1&2\\ 0&1\end{array}\right)$} $\tiny\left(\begin{array}[]{cc|ccc}-1&0&1&1&1\\ 0&-1&1&1&1\\ \hline\cr 0&0&1&0&0\\ 0&0&0&1&0\\ 0&0&0&0&1\end{array}\right)$} $\tiny\left(\begin{array}[]{cc}-1&0\\ 4&1\end{array}\right)$} $\tiny\left(\begin{array}[]{cc|ccc}-1&0&0&0&0\\ 0&-1&0&0&0\\ \hline\cr 2&2&1&0&0\\ 2&2&0&1&0\\ 2&2&0&0&1\end{array}\right)$}
$t_{2}$ $\tiny\left(\begin{array}[]{cc}11&-2\\ 6&-1\end{array}\right)$} $\tiny\left(\begin{array}[]{cc|ccc}5&6&-1&-1&-1\\ 6&5&-1&-1&-1\\ \hline\cr 2&2&-1&0&0\\ 2&2&0&-1&0\\ 2&2&0&0&-1\end{array}\right)$} $\tiny\left(\begin{array}[]{cc}-1&-3\\ 4&11\end{array}\right)$} $\tiny\left(\begin{array}[]{cc|ccc}-1&0&-1&-1&-1\\ 0&-1&-1&-1&-1\\ \hline\cr 2&2&3&4&4\\ 2&2&4&3&4\\ 2&2&4&4&3\end{array}\right)$}
$t_{3}$ $\tiny\left(\begin{array}[]{cc}-11&20\\ -6&11\end{array}\right)$} $\tiny\left(\begin{array}[]{cc|ccc}-5&-6&10&10&10\\ -6&-5&10&10&10\\ \hline\cr-2&-2&3&4&4\\ -2&-2&4&3&4\\ -2&-2&4&4&3\end{array}\right)$} $\tiny\left(\begin{array}[]{cc}-11&-3\\ 40&11\end{array}\right)$} $\tiny\left(\begin{array}[]{cc|ccc}-5&-6&-1&-1&-1\\ -6&-5&-1&-1&-1\\ \hline\cr 20&20&3&4&4\\ 20&20&4&3&4\\ 20&20&4&4&3\end{array}\right)$}
Table 2. $C$- and $G$-matrices for Case (2).